An acid mixture is made by mixing 40% and 70% acid solution. How much of each should be used to get 12 liters of a 50% acid solution?
Answer by mananth(16946)

percent ----------------quantity
40.00% solution-------------x
70.00% solution------------12- x
Total 50.00% ------------12

4000x+7000(12-x)=5000*12
4000x+84000-7000x=60000
4000x-7000x=60000-84000
-3000x=-24000
/-3000
x=8 liters---- 70%
Balance 4 liters 40%